OpenStudy (anonymous):

25 men complete a work in 20 days 15 men leave after some days and the remaining work is completed in 37 1/2 days after how many days did the men leave

ganeshie8 (ganeshie8):

25 men can complete a work in 20 days that means number of `man days` required = 25x20

ganeshie8 (ganeshie8):

say those 15 men left ater \(\large n\) days : \[\large 25\times n + 10\times 37.5 = 25\times 20\] solve \(\large n\)